Four closed surfaces are shown, each with circular top and bottom faces and curved sides. The areas Atop and Abot of the top3.1.10 Select one: a >b>c>d b>a>c>d b>c>ard a = b > c>d c>b> ad
Answer #1

The magnetic flux through a surface is ɸ = A B Cos θ

Where θ is the angle between the normal to the surface and magnetic field. θ= 00 in all the given cases.

For surface A:

Total flux, ɸ = 3 x 8 + 6 x 2 = 30 Wb

For surface B:

Total flux, ɸ = 3 x 2+ 7 x 8 = 62 Wb

For surface C:

Total flux, ɸ = 2 x 9 + 2 x 14 = 46 W

For surface D:

Total flux, ɸ = 2 x 6 - 4 x 2 = 4 Wb

The ranking of the magnitude of magnetic flux ( from greatest to lowest)

B > C > A > D